OpenStudy (anonymous):

Using the formula SA = s2 + 2 × s × l, find the surface area of a square based pyramid with a base length of 3 inches, and a slant height of 7 inches

OpenStudy (ganpat):

substitute s = 3 and l = 7 in given equation..

OpenStudy (anonymous):

So that would make the answer 48?

OpenStudy (ganpat):

SA = s2 + 2 × s × l SA = 3 sq + 2 (3) * 7 SA = 9 + 42 SA = 51.. missed by little.. :)..

OpenStudy (anonymous):

Rats. I see though where I miscalulated. Thank you!